### v1

# DiaSynth - Synthetic Dialogue Generation Framework

Abstract
The scarcity of domain-specific dialogue datasets across various domains, from academic topics to everyday conversations, limits the development of dialogue systems for various applications. Existing research is often constrained either by dialogue datasets that are too general or by niche domain dialogue datasets whose scale does not match the required scale for training dialogue systems. To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains. Our approach differs from existing frameworks by dynamically generating dialogues that incorporate simulated personas, subtopics, and diverse conversational characteristics, using a Large Language Model (LLM) with Chain of Thought (CoT) reasoning to create contextually rich, domain-specific dialogues that closely mimic natural human interactions. DiaSynth produces tailored dialogues that emulate realistic conversations. We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum. The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47%, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the distribution of the in-domain data. The quality of the data generated also scales with the size of LLMs. These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods.
1 Introduction
The development of dialogue systems is an important area of research in the field of natural language processing, due to its wide range of applications ranging from customer service chatbots to virtual personal assistants. The effectiveness of these systems rely heavily on the availability of high-quality, large-scale, domain-specific dialogue datasets. Unfortunately, the lack of high-quality, large-scale datasets across various domains such as academic discussions, healthcare consultations and everyday conversations, poses a big challenge. This scarcity limits the ability of developing dialogue systems that can generalize well across diverse settings and domains.
Existing research has attempted to address this issue by collecting domain specific and real life conversations but they fall into one of the two problematic categories. On the one hand, the conversations in a domain specific dataset are superficial and do not go deep into the domain. On the other hand, niche domain dialogue datasets, while contextually rich, often suffer from limited scale. This imbalance does not allow us to develop dialogue systems for underrepresented domains, where collecting large-scale, high-quality data is either costly or logistically complex.
To address these problems, we introduce DiaSynth, a synthetic dialogue generation framework that produces contextually rich and realistic dialogues tailored to specific domains. DiaSynth, with a Large Language Model (LLM) as the backbone, produces high-quality conversations by simulating personas (representing the characteristics and perspectives of a person) and characteristics, which emulates the implicit qualities of a conversation, such as tone, formality, and agreement levels. By using an LLM as the backbone and Chain of Thoughts (CoT) Wei et al. (2024) as the reasoning mechanism, DiaSynth dynamically generates dialogues that closely mimic real-world conversations, allowing for the rapid generation of tailored dialogue data across a wide range of domains. This approach not only addresses the issue of lack of domain specific dialogue data but also provides a scalable, cost-effective alternative to traditional data collection methods.
To validate the effectiveness of DiaSynth, we evaluated the framework on two criteria: the quality of the generated data and the usability of this data for downstream tasks. The results for the quality criterion showed that the data quality improved with the scale of the model. For the usability criterion, we tested the framework on dialogue summarization task. The results demonstrate that models fine-tuned on DiaSynth-generated data outperformed their base versions by an average of 16.47% across multiple evaluation metrics. Additionally, our comparative analysis reveals that synthetic data captures approximately 90.48% of the performance achieved when fine-tuning on in-domain data, underscoring the framework’s potential as a strong alternate when domain-specific data is unavailable.
The remainder of this paper is organized as follows: Section 2 reviews related work on dialogue datasets and synthetic data generation techniques. Section 3 details the DiaSynth framework, including its architecture and data generation methodology. Section 4 presents our experimental setup, datasets, and evaluation metrics. Section 5 discusses the results of our experiments, comparing the performance of models fine-tuned on DiaSynth-generated data versus in-domain data. Section 6 discusses the limitations of DiaSynth and potential future directions for this research. Finally, Section 7 concludes with a summary of our findings.
2 Related Work
2.1 Personality in Synthetic Data Generation
In recent years, there has been a significant increase in research focused on synthetic dialogue generation, largely driven by advancements in Large Language Models (LLMs). To generate realistic and diverse synthetic data, researchers have incorporated personalities, profiles, and character information when prompting LLMs to generate dialogues Han et al. (2024). By enhancing dialogue realism through the simulation of various personality profiles, utilizing the Big Five personality model, and employing structured prompts, this approach has improved task performance in models fine-tuned on these generated dialogues compared to those trained on general chit-chat datasets.
Moreover, integrating personas into synthetic data generation prompts Chan et al. (2024) has demonstrated that models fine-tuned on personalized synthetic data outperform some LLMs of much larger scales. The inclusion of personas in prompts provides diversity in difficulty levels and ranges within the synthetic data, enabling the models to handle situations of varying complexity.
Our approach involves persona extraction after generating subtopics related to the general topics. This enables the generated dialogues to be more in-depth and specific to those subtopics, enhancing both the scale and quality of domain-specific dialogue generation.
2.2 Prompting Task-Oriented Dialogue Generation
Prompt-based techniques have also emerged as powerful methods for generating high-quality synthetic dialogues, particularly for task-oriented dialogue systems. Steindl et al. (2023) explore the generation of synthetic dialogues from structured prompts, focusing on enhancing task-oriented dialogue systems. Their work demonstrates that prompt engineering can produce dialogues that are contextually appropriate and improve system performance by aligning synthetic data more closely with real-world requirements.
To achieve a higher quantity, diversity, and creativity in human-written instruction data, Wang et al. (2022) propose inputting prompts to LLMs to generate instructions based on a small set of seed human-written instructions. This approach aligns the expanded training data more closely with desired task objectives and allows for iterative improvements, producing more nuanced and effective dialogues that meet specific task demands.
Similarly, our study expands topics into subtopics, ensuring that the generated dialogues provide more in-depth and high-quality conversations. By doing so, we aim to produce synthetic data that not only covers a broader range of scenarios but also delves deeper into each topic, thereby enhancing the overall effectiveness of the dialogue systems trained on this data.
3 DiaSynth
DiaSynth is a synthetic dialogue generation framework designed to address the scarcity of high-quality, large-scale, domain-specific dialogue datasets. The framework functions with an LLM as the backbone and CoT as the reasoning mechanism, to simulate diverse real world dialogues that capture the complexity and nuances of human conversations.
DiaSynth operates by taking in a list of topics from the user based on which the conversations should be generated. The users can optionally provide few-shot examples of the format in which they want the dialogue to be generated. Dialogues that are generated directly from the given list of topics will be too superficial since the topics given by the user will not be specific. To overcome this lack of specificity, we generate $m$ sub topics for each of the $n$ topics given by the user. Generating dialogues from the subtopics will have specificity but the dialogues will lack variety. This is because every dialogue is influenced implicitly by the personas of the people involved in the dialogue and, other characteristics such as the location, emotion and more. To increase both variety and depth in the dialogues generated, we generate $p$ personas for each of the $m$ subtopics. The dialogues are then generated for all possible combinations of personas belonging to each of the subtopics. To further ground the dialogues in various settings and characteristics, we employ CoT reasoning during the generation process. DiaSynth employs CoT to reason about the settings and characteristics of a dialogue, as listed in Table 1, ensuring that the dialogues are contextually rich and realistic. This three-stage pipeline not only guarantees the quality of the generated dialogues but also allows for exponential scalability, as illustrated programmatically in Appendix A.
3.1 Subtopic Generation
Subtopic generation is a crucial step in DiaSynth’s pipeline, since it enhances the specificity and depth of the dialogues that will be generated later. For each primary topic given by the user, DiaSynth generates multiple subtopics, effectively narrowing the focus of the conversation. This breakdown is necessary because the primary topics are often too general to generate contextually rich dialogues on their own. For instance, a topic like “healthcare” can be expanded into subtopics such as “doctor-patient consultations,” “mental health discussions,” and “medical diagnostics,” each of which offers a more focused context for dialogue generation. To achieve this, DiaSynth prompts an LLM to generate the user specified number of subtopics for each primary topic. We also run a similarity check between each of the subtopics and remove subtopics that are too similar to other subtopics using a threshold.
3.2 Persona Generation
Personas of the individuals involved in a conversation are primary influencers in determining how a conversation pans out. Using random personas from persona datasets and prompting the LLM to simulate a dialogue between them about a random topic often leads to superficial dialogues that lack depth and contextual richness. To address this issue, DiaSynth generates a user-specified number of personas for each subtopic, ensuring that the personas are conditioned on the subtopic context. This conditioning prompts the LLM to create personas that are most likely to engage in a meaningful dialogue about the subtopic, such as a medical professional and a patient discussing "medical diagnostics" or a researcher and a student talking about "academic publishing." We also run a similarity check for the personas too. The conditioned persona generation is crucial because it ensures that future dialogues will not only be contextually rich but also exhibit a high level of depth. Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting. Notably, during our experiments, the similarity checks indicated no significant overlap among the dialogues, suggesting that the inherent diversity of the dialogues stems from DiaSynth’s multi-stage generation pipeline.
Figure 1: Overview of the Dialogue Synthesis Framework: The diagram shows the step-by-step process of generating dialogues from user-provided topics, involving sub-topic generation, persona generation, dialogue generation, and filtering based on calculated metrics for downstream tasks.
3.3 Dialogue Generation
The final stage in DiaSynth’s pipeline is the generation of dialogues, where all the components—subtopics, personas, and characteristics—converge to create contextually rich and realistic conversations. This step uses an LLM as the backbone and CoT as the reasoning mechanism, allowing the model to simulate dialogues that incorporate various aspects of human interaction. In this stage, for each subtopic, DiaSynth generates dialogues by pairing every possible combination of personas within that subtopic. The generation process also takes into account the predefined dialogue characteristics as listed in Table 1—such as emotional state, formality level, and familiarity between speakers—to guide the conversation flow and style. These characteristics are explicitly defined in the CoT prompt, allowing the LLM to generate dialogues that are both contextually appropriate and realistic.
Characteristic Description
Age and Gender Defines the demographic details of both personas, influencing the style, tone, and cultural references in the dialogue.
Familiarity Level Indicates how well the speakers know each other, which affects the formality, trust, and depth of the conversation.
Emotional States Specifies the emotional condition of the speakers (e.g., angry, happy, sad), impacting the tone and progression of the dialogue.
Formality Level Determines the level of politeness or casualness in the dialogue, which depends on context and familiarity between speakers.
Duration of the Conversation Suggests the intended length of the conversation, which can influence the dialogue’s depth and complexity.
Communication Medium Defines the medium of the conversation (e.g., face-to-face, phone, text), which affects the dialogue style and expressions used.
Topic of the Conversation Provides the main subject around which the dialogue is centered, guiding the content and direction of the conversation.
Location of the Conversation Indicates where the dialogue takes place, adding context that can influence the dialogue’s formality and content.
Agreement or Disagreement Shows whether the speakers agree or disagree on the topic, which drives the dialogue dynamics and interaction style.
Natural Dialogue Features Incorporates elements like fillers (e.g., "umm," "uh"), pauses, and slang, adding authenticity and naturalness to the dialogue.
Table 1: Characteristics of the Dialogue for CoT Prompt
4 Experimental Setup
In this section, we detail the experimental setup used to evaluate the effectiveness of DiaSynth. Our evaluation focuses on two criteria - quality of the dialogues generated and usability of the dialogues generated for a downstream task. Quality of the dialogues is evaluated using metrics such as FED, GPTScore, and G-Eval. The usability of the dialogues generated for a downstream task is evaluated by choosing summarization as the downstream task.
4.1 Quality of the dialogues
To evaluate the quality of the dialogues, we employ the metrics that have been developed for evaluating the quality of text generated by LLMs. The metrics are as follows,
  • FED Mehri and Eskenazi (2020) - FED evaluates the quality of dialogue by appending a positive and a negative utterance to the end of the dialogue and measuring the difference in probabilities assigned to the positive and negative utterances.
  • GPTScore Fu et al. (2024) - GPTScore assesses dialogue quality based on various characteristics such as coherence, diversity, and others. The evaluation involves asking an LLM whether a dialogue meets a specific criterion. The score for each criterion is determined by the probability of the LLM responding affirmatively (’Yes’).
  • G-Eval Liu et al. (2023) - G-Eval evaluates dialogue quality by prompting an LLM to rate the dialogue on a scale of 1 to 3 across multiple criteria. The final score is calculated using a weighted average of each rating, with weights derived from the probability distribution of the LLM’s assigned ratings.
To validate the framework across models and also domains, we generate data using three open source LLMs, one closed source LLM and also use few shot examples from two different dialogue datasets. The four LLMs used are - Phi-3 (phi-3-mini-128k-instruct (3.8B)  [^1] Phi-3 HF link) Abdin et al. (2024), InternLM-2.5 (internlm2_5-7b-chat (7.74B) [^2] InternLM-2.5 HF link) Cai et al. (2024), LLaMA-3 (llama3-8b-instruct (8.03B) [^3] LLaMA-3 HF link) Dubey et al. (2024) and GPT-4o. The two different dialogue datasets that were used are DialogueSum Chen et al. (2021) and SAMSum Gliwa et al. (2019)
4.2 Downstream Task - Summarization
To evaluate the usability of the dialogues generated by DiaSynth, we choose summarization as the downstream task. Summarization is an essential application of dialogue systems, where the goal is to generate concise, and contextually relevant summaries. We use three established evaluation metrics—QAGS, BERTScore, and ROUGE-L—to assess the performance of summarization models fine-tuned on DiaSynth-generated data.
  • QAGS Wang et al. (2020) - QAGS (Question Answering Guided Summarization) evaluates the factual consistency of generated summaries by automatically generating questions based on the summary and then comparing the answers obtained from the summary to those from the source dialogue. In our work, instead of generating questions from the summaries, we generate them from the source dialogues, allowing us to directly compare the consistency and accuracy of summaries generated from different datasets containing the same dialogues.
  • BERTScore Zhang* et al. (2020) - BERTScore measures the semantic similarity between the generated summary and the reference summary using contextual embeddings.
  • ROUGE-L Lin (2004) - ROUGE-L is a widely used evaluation metric that measures the overlap of longest common subsequences (LCS) between the generated summary and the reference summary.
We fine-tune pretrained models that perform well on summarization such as DistilBART, BART Lewis et al. (2020), T5 Raffel et al. (2020) and LED Beltagy et al. (2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics. We evaluate the effectiveness of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
5 Results
This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs as the backbone and varying few-shot examples. Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets. These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples. In total, eight distinct datasets were generated using DiaSynth by pairing each LLM backbone with the two sets of few-shot examples, resulting in all possible combinations. For each dataset, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic. The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 2.
LLM Few-Shot Number of Samples Avg. number of turns Avg. number of tokens per turn Diversity (rouge-l)
Phi-3 DialogueSum 1215 9.13 20.38 0.27
Phi-3 SAMSum 1410 13.98 13.94 0.27
InternLM-2.5 DialogueSum 1035 9.23 27.98 0.30
InternLM-2.5 SAMSum 1135 13.96 19.07 0.29
LLaMA-3 DialogueSum 1154 6.86 31.99 0.29
LLaMA-3 SAMSum 1195 10.54 20.41 0.29
GPT-4o DialogSum 1375 15.16 15.96 0.29
GPT-4o SAMSum 1380 15.43 13.53 to 0.28
Table 2: Data Statistics
5.1 Quality of the Dialogues
The quality of the synthetic datasets produced by DiaSynth was evaluated using FED, GPTScore, and G-Eval metrics, as detailed in Tables 3, 4, and 5. The results reveal distinct variations in performance across different model and dataset configurations, reflecting the unique characteristics of each.
The FED scores in Table 3 indicate that LLaMA-3 and GPT-4o consistently outperform Phi-3 and InternLM-2.5, particularly in coherence, consistency, and diversity, with LLaMA-3 leading in the DialogSum dataset (e.g., coherence 0.9684, consistency 0.9570, diversity 0.9596) and GPT-4o excelling in the SAMSum dataset, especially in coherence (0.9939) and consistency (0.9878). Table 4 further supports LLaMA-3’s adaptability and error-handling abilities, as it outperforms others in diversity (0.0430) and inquisitiveness (0.0415) in DialogSum, while maintaining an edge in SAMSum for diversity (0.0655) and inquisitiveness (0.0622). Though GPT-4o remains strong in coherence and consistency, Phi-3 shows competitive coherence in SAMSum (0.0325). Finally, Table 5 highlights GPT-4o’s dominance in engagingness and naturalness with perfect scores (3.0) for DialogSum, while InternLM-2.5 stands out in coherence (2.9990) and groundedness (2.9973) for DialogSum, and coherence (2.9983) and groundedness (2.9952) for SAMSum, suggesting it maintains high factual accuracy. The contrasting performance of GPT-4o on the DialogSum and SAMSum datasets can be attributed to the differing structures of the dialogues in these datasets. DialogSum consists of more formal and structured dialogues, which aligns with the typical response style of GPT-4o, leading to its stronger performance. In contrast, SAMSum contains more casual, human-like conversations, which might explain GPT-4o’s relatively poorer performance, as it may not adapt as well to the informal, spontaneous nature of such dialogues. Overall, while GPT-4o excels in natural and engaging dialogue, LLaMA-3 offers the most versatility, and InternLM-2.5 provides a strong alternative with high coherence and groundedness.
The observed superiority of LLaMA-3 over GPT-4o is surprising because an 8 billion 8-bit quantized model not only competes with but also performs better than GPT-4o in certain metrics. We hypothesize that this could be due to the way GPT-4o was trained, which might make it more constrained in its responses, whereas LLaMA-3, being an open-source model, operates with fewer restrictions. This allows LLaMA-3 to exhibit greater flexibility, diversity, and adaptability in generating dialogues, potentially explaining its better performance in certain metrics. These characteristics can be seen in criteria like ’inquisitiveness’ and ’likeability’ in Table 4 and, ’depth’ and ’diverse’ in Table 3. These results suggest that for building human-like data generation frameworks, open-source LLMs are a more suitable choice than closed-source LLMs. The minimal constraints on response formatting during the training of open-source models enable them to generate more diverse, flexible, and human-like dialogues, making them better suited for tasks requiring natural and conversational interactions.
coherent error recovery consistent diverse depth likeable understand flexible informative inquisitive
DIALOGUESUM
Phi-3 0.9536 0.9440 0.9540 0.9534 0.9521 -0.0005 0.9353 -3.96E-05 0.0009 -0.0033
InternLM-2.5 0.8439 0.8313 0.8359 0.8353 0.8352 0.0048 0.8278 -0.0046 0.0042 0.0069
LLaMA-3 0.9684 0.9522 0.9570 0.9596 0.9592 0.0032 0.9453 -0.0063 0.0063 0.0105
GPT-4o 0.9525 0.9407 0.9417 0.9423 0.9425 0.0121 0.9368 -0.0027 0.0085 0.0144
SAMSUM
Phi-3 0.9161 0.9088 0.9199 0.9161 0.9130 -0.0004 0.9014 -0.0024 0.0034 -0.0040
InternLM-2.5 0.8746 0.8647 0.8734 0.8655 0.8661 0.0033 0.8582 -0.0019 0.0106 0.0028
LLaMA-3 0.9829 0.9677 0.9757 0.9712 0.9731 0.0003 0.9593 -0.0048 0.0100 0.0029
GPT-4o 0.9939 0.9876 0.9878 0.9836 0.9824 0.0083 0.9788 0.0004 0.0141 0.006
Table 3: FED scores
coherence diversity flexibility understandability inquisitiveness consistency informativeness likeability depth error recovery
DIALOGUESUM
Phi-3 0.0286 0.0310 0.0218 0.0193 0.0363 0.0369 0.0172 0.0213 0.0117 0.0342
InternLM-2.5 0.0069 0.0196 0.0084 0.0061 0.0244 0.0137 0.0148 0.0050 0.0080 0.0197
LLaMA-3 0.0189 0.0430 0.0186 0.0220 0.0415 0.0321 0.0318 0.0110 0.0201 0.0440
GPT-4o 0.0039 0.0156 0.0059 0.0039 0.018 0.0080 0.0097 0.0026 0.0053 0.0135
SAMSUM
Phi-3 0.0325 0.0372 0.0260 0.0270 0.0395 0.0415 0.0201 0.0232 0.0126 0.0290
InternLM-2.5 0.0128 0.0408 0.0194 0.0174 0.0504 0.0306 0.0328 0.0142 0.0177 0.0407
LLaMA-3 0.0288 0.0655 0.0306 0.0365 0.0622 0.0612 0.0542 0.0168 0.0270 0.0558
GPT-4o 0.0055 0.0162 0.0094 0.0084 0.0186 0.0119 0.0129 0.0038 0.0079 0.0199
Table 4: GPTScore
engagingness naturalness coherence groundedness
DIALOGUESUM
Phi-3 2.5236 2.7238 2.6308 2.5557
InternLM-2.5 2.9995 2.9989 2.9990 2.9973
LLaMA-3 2.9987 2.9988 2.9972 2.9935
GPT-4o 3 3 3 2.9975
SAMSUM
Phi-3 2.4623 2.6821 2.5848 2.5060
InternLM-2.5 2.9992 2.9969 2.9983 2.9952
LLaMA-3 2.9976 2.9971 2.9969 2.9916
GPT-4o 2.9994 2.9977 2.9982 2.9944
Table 5: G_EVAL
5.2 Fine-tuning and Performance Results
To validate the usability of the synthetic data generated using DiaSynth, we fine-tuned and evaluated several pretrained language models on the task of dialogue summarization. The summaries for dialogues generated by different LLMs were created using the corresponding LLMs through prompting. The pretrained models used for evaluation include DistilBART, BART, T5, and LED.
The experimental setup is designed as follows:
  • All reported metrics are based on the validation and test sets of DialogSum and SAMSum to maintain consistency in evaluation.
  • To assess the effectiveness of DiaSynth-generated data, we compare the performance of models fine-tuned on DiaSynth data with their corresponding base models, which are evaluated without any fine-tuning.
  • Given that the training sets of DialogSum and SAMSum are larger than the synthetic datasets generated in this study, we randomly sampled instances from the in-domain training sets to match the size of the DiaSynth-generated data. This ensures a fair comparison in terms of training scale.
  • The primary goal of the experiment is to quantify the percentage improvement offered by models fine-tuned on DiaSynth-generated data compared to their base versions and to evaluate the extent to which the generated data capture the distributional characteristics of the in-domain data.
The results presented in Table 6 highlight the performance of each model before and after fine-tuning. Notably, base models such as DistilBART and LED exhibit relatively high QAGS scores before fine-tuning, primarily because these models tend to repetitively generate content similar to the input dialogues, thereby artificially inflating the factual consistency metric. This repetition, while beneficial for preserving factual alignment, does not necessarily translate into improved quality or informativeness of the summaries.
After fine-tuning on DiaSynth-generated data, models show a significant increase in BERTScore and ROUGE-L compared to their base versions, demonstrating DiaSynth’s effectiveness in enhancing model performance for downstream tasks. This result suggests that DiaSynth is valuable for generating dialogue data in domains with limited training resources. Notably, models fine-tuned on GPT-4o-generated data performed best for DialogSum, while those fine-tuned on LLaMA-3-generated data excelled on SAMSum, further reinforcing that open-source LLMs are more adept at generating informal, casual dialogues like those in SAMSum. The percentage improvement of models fine-tuned on LLaMA-3-generated data was calculated using Equation 1, showing an average improvement of 16.47% across the 24 reported results. We chose to calculate on LLaMA-3 because across both quality and usability it performs better than GPT-4o.
Percentage Improvement=Score After Fine-tuningScore Before Fine-tuningScore Before $\text{Percentage Improvement}=\frac{\text{Score After Fine-tuning}-\text{ScoreBefore Fine-tuning}}{\text{Score Before Fine-tuning}}\times 100$ (1)
Coverage Percentage=Finetuned Score on DiaSynth DataFinetuned Score on In-Domain Base $\text{Coverage Percentage}=\frac{\text{Finetuned Score on DiaSynth Data}}{\text{Finetuned Score on In-Domain Base Data}}\times 100$ (2)
To assess the percentage coverage of the distributional characteristics of the in-domain data by the synthetically generated data, we use Equation 2 and the scores of models fine-tuned on LLaMA-3 generated data. Across the 24 reported results, the overall coverage percentage of the LLaMA-3 generated data is 90.48%. Notably, the QAGS scores of models fine-tuned on synthetic data surpass those of models trained on in-domain data, suggesting that synthetic data can match or even exceed in-domain data performance in some aspects. Excluding QAGS, the coverage percentage is calculated to be 77.07%, highlighting that even with moderate LLMs of small scale (3B - 8B), high-quality synthetic dialogue datasets can be effectively created.
Model Before Finetuning Finetuning on Base Data
QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L
DIALOGSUM
distillbart-cnn 0.6134 0.5093 0.1950 0.5586 0.7005 0.3367
bart-base 0.7007 0.5274 0.1375 0.4789 0.6868 0.2969
t5-base 0.5901 0.5491 0.1812 0.4766 0.6953 0.2986
led-base-16384 0.8261 0.5471 0.1634 0.4872 0.7084 0.3165
SAMSUM
distillbart-cnn 0.6627 0.5500 0.2394 0.6041 0.6849 0.3578
bart-base 0.7563 0.4389 0.1765 0.5302 0.6520 0.3049
t5-base 0.5574 0.4190 0.1237 0.5460 0.6448 0.3000
led-base-16384 0.7429 0.4310 0.1812 0.5440 0.6522 0.3175
Table 6: Performance of models before and after finetuning on in-domain data
Model Phi-3 InternLM-2.5 LLaMA-3 GPT-4o
QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L QAGS BERTScore ROUGE-L
DIALOGUESUM
distillbart-cnn 0.6588 0.5778 0.2187 0.6420 0.6008 0.2167 0.6586 0.6161 0.2040 0.6713 0.6242 0.2014
bart-base 0.5355 0.5958 0.2029 0.5418 0.6212 0.1897 0.5825 0.6033 0.1789 0.5590 0.6039 0.1769
t5-base 0.5937 0.5949 0.2047 0.5825 0.5941 0.1878 0.6034 0.6172 0.1959 0.6305 0.6319 0.2044
led-base-16384 0.5358 0.6129 0.2109 0.5189 0.6027 0.1606 0.5697 0.6302 0.1999 0.5791 0.6308 0.1989
SAMSUM
distillbart-cnn 0.6585 0.5931 0.2262 0.6388 0.6066 0.2422 0.6849 0.6029 0.2374 0.6757 0.6029 0.2291
bart-base 0.5648 0.5665 0.2146 0.5435 0.5663 0.2021 0.6132 0.5899 0.2345 0.5707 0.5808 0.2154
t5-base 0.5905 0.5397 0.2085 0.5457 0.5193 0.1854 0.6412 0.5054 0.1976 0.6023 0.5419 0.1979
led-base-16384 0.5883 0.5477 0.2289 0.5457 0.5615 0.2167 0.5917 0.5785 0.2390 0.5738 0.569 0.2298
Table 7: Performance metrics
6 Limitations
Despite the promising results, our approach has some limitations. Firstly, different LLMs exhibit varied performance based on the dialogue structure, with certain models like LLaMA-3 performing better for more informal dialogues (e.g., SAMSum) and others like GPT-4o excelling in structured, formal dialogues (e.g., DialogSum). This indicates that there is no single model that can universally handle all types of dialogue structures, requiring a tailored choice of LLM depending on the target application.
Secondly, the generation process may suffer from a lack of knowledge on certain topics, especially in cases where the LLMs were not sufficiently trained on those domains. Additionally, our framework relies on zero-shot generation for personas and sub-topics, which, while flexible, can sometimes result in less coherent or less accurate persona simulation, as it is not fine-tuned for specific contexts. These limitations suggest directions for future work, such as combining LLMs to leverage their strengths or incorporating more topic-specific training to enhance knowledge coverage.
7 Conclusion
In this paper, we introduced DiaSynth, a synthetic dialogue generation framework capable of producing high-quality, contextually rich dialogues across a wide range of domains. Our experiments demonstrated that models fine-tuned on DiaSynth-generated data exhibit significant improvements in downstream tasks, as evidenced by substantial increases in BERTScore and ROUGE-L compared to their base models. These results highlight the potential of DiaSynth as an effective tool for generating dialogue data, particularly in domains where training data is scarce.
Furthermore, our analysis showed that different LLMs excel in different dialogue structures, with LLaMA-3 performing better for informal dialogues and GPT-4o for more structured settings. This insight suggests that leveraging open-source LLMs may be more advantageous for generating human-like conversational data. Despite certain limitations, such as varying LLM performance across dialogue types and knowledge gaps in zero-shot generation, DiaSynth presents a promising approach to dialogue data generation and offers a valuable resource for future advancements in building more sophisticated and adaptable dialogue systems.
Appendix A Scalability of DiaSynth
This section illustrates the scalability of DiaSynth with a python program and examples.
1def calculate_total_dialogs(n, m, p):
2 """
3 Calculate the total number of dialogs generated.
4
5 Parameters:
6 n (int): Number of topics
7 m (int): Number of subtopics per topic
8 p (int): Number of personas per subtopic
9
10 Returns:
11 int: Total number of dialogs generated
12 """
13 # Calculate the total number of subtopics
14 total_subtopics = n * m
15
16 # Calculate the number of dialogs for each subtopic using combinations of personas
17 dialogs_per_subtopic = (p * (p - 1)) // 2
18
19 # Total dialogs generated across all subtopics
20 total_dialogs = total_subtopics * dialogs_per_subtopic
21
22 return total_dialogs
Listing 1: Calculation of Total Dialogs Generated
A.1 Example 1
  • Number of topics ($n$): 10
  • Number of subtopics per topic ($m$): 5
  • Number of personas per subtopic ($p$): 3
Calculation:
Total subtopics $=n\times m=10\times 5=50$
Dialogues per subtopic $=\frac{p\times(p-1)}{2}=\frac{3\times 2}{2}=3$
Total dialogues $=50\times 3=150$
This setup generates 150 dialogues.
A.2 Example 2
  • Number of topics ($n$): 20
  • Number of subtopics per topic ($m$): 4
  • Number of personas per subtopic ($p$): 5
Calculation:
Total subtopics $=20\times 4=80$
Dialogues per subtopic $=\frac{5\times 4}{2}=10$
Total dialogues $=80\times 10=800$
This setup generates 800 dialogues.
A.3 Example 3
  • Number of topics ($n$): 15
  • Number of subtopics per topic ($m$): 6
  • Number of personas per subtopic ($p$): 10
Calculation:
Total subtopics $=15\times 6=90$
Dialogues per subtopic $=\frac{10\times 9}{2}=45$
Total dialogues $=90\times 45=4050$
This setup generates 4050 dialogues.
A.4 Scaling Observations
  • Linear Scaling with Topics and Subtopics: Increasing the number of topics or subtopics results in a linear increase in the total number of dialogues, making it straightforward to expand the scope of dialogue generation.
  • Exponential Scaling with Personas: The number of dialogues scales exponentially as the number of personas increases because each additional persona allows for more combinations, making the framework highly scalable for complex scenarios.
  • Practical Use Case: For specific domains like academic, healthcare, or business, these parameters can be adjusted to generate thousands of dialogues to fit the needs of various applications such as training chatbots, virtual assistants, or dialogue-based assessments.
These examples illustrate DiaSynth’s potential for rapid and scalable generation of dialogues, which can be tailored to different domains by simply adjusting the input parameters.